# A lower bound on the right-handed neutrino mass from wash-in leptogenesis

### v1

@@ Abstract
Leptogenesis is an attractive scenario for the generation of the baryon asymmetry of the Universe that relies on the dynamics of right-handed neutrinos (RHNs) in the seesaw extension of the Standard Model. In standard thermal leptogenesis, the RHN mass scale $M_{N}$ is subject to the Davidson–Ibarra bound, $M_{N}\gtrsim 10^{9}\,\textrm{GeV}$, which builds on the assumption that RHN decays are responsible for the violation of both charge-parity invariance ($CP$) and baryon-minus-lepton number ($B\!-\!L$). In this paper, we relax this assumption in the context of the more general framework of wash-in leptogenesis, in which $CP$ violation is encoded in the initial conditions and the only remaining task of the RHN decays is to violate $B\!-\!L$. Solving the relevant set of Boltzmann equations, we find that, in wash-in leptogenesis, the RHN mass scale can be as low as 7 TeV.
preprint: MITP-25-002preprint: MS-TP-24-01
@@
Introduction — The baryon asymmetry of the Universe (BAU), commonly quantified in terms of the baryon-to-photon ratio $\eta_{B}^{\rm obs}=n_{B}^{0}/n_{\gamma}^{0}\simeq 6.12\times 10^{-10}$ Planck:2018vyg ; ParticleDataGroup:2024cfk , cannot be explained by the Standard Model (SM) of particle physics and hence provides evidence for new physics beyond the SM (BSM) Sakharov:1967dj . An attractive BSM scenario capable of dynamically generating the BAU in the early Universe consists of baryogenesis via leptogenesis Fukugita:1986hr , which relies on the interactions of right-handed neutrinos (RHNs) in the seesaw extension of the SM Minkowski:1977sc ; Yanagida:1979as ; Yanagida:1980xy ; Gell-Mann:1979vob ; Mohapatra:1979ia , thus linking the generation of the BAU to the origin of the tiny masses of the left-handed neutrinos in the SM Buchmuller:2005eh ; Chun:2017spz ; Bodeker:2020ghk . In its standard implementation, leptogenesis proceeds via the out-of-equilibrium decays of heavy RHNs, which simultaneously violate charge–parity invariance ($CP$) and baryon-minus-lepton number ($B\!-\!L$). These decays first produce a primordial lepton asymmetry, which is subsequently converted by means of the chemical transport in the thermal plasma, including electroweak sphalerons Kuzmin:1985mm , to a primordial baryon asymmetry.
@@
The requirement of sufficient $CP$ violation in RHN decays bounds the RHN mass scale from below, $M_{N}\gtrsim 10^{9}\,\textrm{GeV}$. This constraint is known as the Davidson–Ibarra bound Davidson:2002qv ; Buchmuller:2002rq (see Ref. Garbrecht:2024xfs for a recent refinement). It assumes a hierarchical RHN mass spectrum (e.g., $M_{3}\gg M_{2}\gg M_{1}\equiv M_{N}$ in models with three RHNs) and can be circumvented in the case of a quasi-degenerate RHN mass spectrum (e.g., $M_{3}\gg M_{2}\simeq M_{1}$ or $M_{3}\simeq M_{2}\simeq M_{1}$), which gives rise to the scenario of resonant leptogenesis Pilaftsis:2003gt ; Pilaftsis:2005rv (see also Refs. Akhmedov:1998qx ; Klaric:2020phc ; Klaric:2021cpi ). In this paper, we will, however, ignore the possibility of resonantly enhanced $CP$ asymmetry in RHN decays and focus instead on an alternative route to low-scale leptogenesis at RHN masses $M_{N}\ll 10^{9}\,\textrm{GeV}$: wash-in leptogenesis Domcke:2020quw .
@@
The key idea behind wash-in leptogenesis is to liberate the RHN decays from the requirement of sufficient $CP$ violation. Indeed, in wash-in leptogenesis, RHN interactions are treated as mere spectator processes and on the same footing as the electroweak sphalerons. From the perspective of wash-in leptogenesis, sphalerons and RHN decays represent nothing but sources of $B\!+\!L$ violation and $B\!-\!L$ violation, respectively. $CP$ violation, on the other hand, is attributed to new $CP$-violating dynamics at high temperatures, $T\gg M_{N}$, which set the required $CP$-violating initial conditions for wash-in leptogenesis. In this sense, wash-in leptogenesis does not represent a complete baryogenesis scenario on its own, but calls for an ultraviolet completion — a chargegenesis mechanism that is responsible for the generation of primordial $CP$-violating charge asymmetries at high energies. Wash-in leptogenesis then acts on the nontrivial configuration of chemical potentials induced by chargegenesis and reshuffles the primordial charge asymmetries in a $B\!-\!L$-violating manner. The SM notably offers more than ten conserved global charges at high temperatures that could be produced during chargegenesis Domcke:2020quw , which renders baryogenesis via leptogenesis via chargegenesis significantly more versatile than baryogenesis via leptogenesis. As a consequence, the range of possible chargegenesis scenarios is extremely broad; primordial charge asymmetries suitable for wash-in leptogenesis can, e.g., be produced during axion inflation Domcke:2022kfs , by evaporating primordial black holes Schmitz:2023pfy , in the decay of heavy Higgs doublets Mukaida:2024eqi , or via new nonminimal gravitational interactions Mojahed:2024yus .
@@
As the temperature of the thermal bath drops, the last global charge to be erased by the SM interactions is the asymmetry between right-handed electrons and left-handed positrons, $q_{e}=n_{e}-\bar{n}_{e}$. Therefore, as long as the temperature scale of wash-in leptogenesis exceeds the equilibration temperature of the electron Yukawa interaction, $T_{e}\simeq 85\,\textrm{TeV}$ Bodeker:2019ajh , it is always possible to realize successful wash-in leptogenesis by assuming a nonzero primordial electron charge, $q_{e}^{\rm ini}\neq 0$. Correspondingly, the authors of Ref. Domcke:2020quw concluded that wash-in leptogenesis can take place at temperatures as low as $T\simeq 100\,\textrm{TeV}$. The purpose of the present paper is to make this statement more precise and study wash-in leptogenesis at $T\lesssim T_{e}$, where $q_{e}/T^{3}$ can no longer be treated as a constant, i.e., in a temperature regime where the asymmetry wash-in by RHN processes competes with the wash-out of the primordial $q_{e}$ input charge due to the progressing equilibration of the electron Yukawa interaction. To this end, we shall extend the treatment in Ref. Domcke:2020quw and supplement the Boltzmann equations for the three flavor asymmetries $q_{\Delta_{\alpha}}$, where $\Delta_{\alpha}=B/3-L_{\alpha}$ and $\alpha=e,\mu,\tau$, that are being produced during wash-in leptogenesis by a Boltzmann equation for the time-dependent electron charge $q_{e}$. Solving this coupled set of Boltzmann equations will then allow us to conclude that wash-in leptogenesis is able to produce the BAU for RHN masses as low as $M_{N}\simeq 7\,\textrm{TeV}$. In summary, in standard thermal leptogenesis where RHN decays act as a source of both $CP$ and $B\!-\!L$ violation, the RHN mass scale must satisfy the Davidson–Ibarra bound, $M_{N}\gtrsim 10^{9}\,\textrm{GeV}$; if RHN decays only act as a source of $B\!-\!L$ violation, it suffices to demand that $M_{N}\gtrsim 7\,\textrm{TeV}$.
@@
$-\left(\partial_{t}+3H\right)q_{\Delta_{\alpha}}=\varepsilon_{1\alpha}\,\Gamma_{1}\left(n_{N_{1}}-n_{N_{1}}^{\rm eq}\right)-\sum_{\beta}\gamma_{\alpha\beta}^{\rm w}\,\frac{\mu_{\ell_{\beta}}+\mu_{\phi}}{T}\,.$ (1)
Here, the negative sign on the left-hand side reflects the fact that the lepton flavor number $L_{\alpha}$ enters $\Delta_{\alpha}$ with a negative sign. The first term on the right-hand side is the source term for $q_{\Delta_{\alpha}}$ in standard thermal leptogenesis. In the following, we will assume vanishingly small $CP$ violation in the RHN sector, $\varepsilon_{1\alpha}\lll 1$, and simply ignore this term. Furthermore, the second term on the right-hand side is the wash-out term in standard thermal leptogenesis, where the total wash-out interaction density receives contributions from RHN inverse decays as well as from lepton-number violating ($\Delta L=2$) and lepton-flavor-violating ($\Delta L=0$) processes. This interaction density is multiplied by $\mu_{\ell_{\beta}}+\mu_{\phi}$, i.e., the sum of the chemical potentials of the SM charged-lepton and Higgs doublets. At $T\sim T_{e}$, the chemical potentials of the 16 SM particle multiplets, $\mu_{i}$ ($i=e,\mu,\tau,\ell_{e},\ell_{\mu},\ell_{\tau},u,c,t,d,s,b,Q_{1},Q_{2},Q_{3},\phi$), are subject to 16 constraints: 11 conditions for 11 linearly independent Yukawa and sphaleron interactions in chemical equilibrium and 5 conservation laws for 5 conserved global charges ($\Delta_{e,\mu,\tau}$, hypercharge $Y$, and the asymmetry in right-handed electrons). Solving this linear system of constraint equations for the 16 chemical potentials $\mu_{i}$, one finds Domcke:2020kcp ; Domcke:2020quw
@@
where the second term on the right-hand side contains the standard flavor coupling matrix Barbieri:1999ma ; Abada:2006fw ; Nardi:2006fx ; Abada:2006ea ; Blanchet:2006be ; Antusch:2006cw induced by the spectator processes in the thermal bath Buchmuller:2001sr ; Garbrecht:2014kda ; Garbrecht:2019zaa , while the first term is the relevant term for wash-in leptogenesis.
In presence of a primordial charge asymmetry, $\bar{\mu}_{e}\neq 0$, the processes contributing to the wash-out rate will not drive the $q_{\Delta_{\alpha}}$ to zero. Quite the opposite, these processes will now wash in nonzero lepton flavor asymmetries $q_{\Delta_{\alpha}}\equiv\nicefrac{{1}}{{6}}\>\bar{\mu}_{\Delta_{\alpha}}T^{2}$ proportional to $q_{e}\equiv\nicefrac{{1}}{{6}}\>\bar{\mu}_{e}T^{2}$, even if $q_{\Delta_{\alpha}}=0$ initially. Assuming $q_{e}/T^{3}=\textrm{const}$, wash-in leptogenesis in the strong wash-in regime (conventionally, strong wash-out regime) at $T\gtrsim T_{e}$ thus results in a $B\!-\!L$ asymmetry $q_{B-L}=-\nicefrac{{3}}{{10}}\,q_{e}$ Domcke:2020quw .
Around $T\sim T_{e}$, the assumption of constant $q_{e}/T^{3}$, however, begins to break down and $q_{e}/T^{3}$ needs to be described as a time-dependent quantity. The relevant Boltzmann equation for $q_{e}$ around the equilibration temperature of the electron Yukawa interaction has been derived in Ref. Bodeker:2019ajh ,
@@
Meanwhile, the electron equilibration rate can be written as $\Gamma_{e}=G_{e}\,T$, where the coefficient $G_{e}$ receives contributions from $2\rightarrow 2$ processes involving one electron Yukawa vertex as well as from $1\leftrightarrow 2$ decays and inverse decays of the SM Higgs bosons, right-handed electrons and left-handed leptons. In fact, the latter processes also receive corrections from $1n\leftrightarrow 2n$ scatterings with soft gauge boson exchanges, which can be taken care of by Landau–Pomeranchuk–Migdal (LPM) resummation Landau:1953um ; Landau:1953gr ; Migdal:1956tc . Taking all these processes into account, the authors of Ref. Bodeker:2019ajh find
@@
where $c_{t}$, $c_{\ell}$, and $c_{e}$ receive contributions from both $2\rightarrow 2$ processes and LPM-resummed $1n\leftrightarrow 2n$ processes,
@@
Four example solutions of our set of Boltzmann equations for these initial conditions are shown in Fig. 1. For a given numerical solution, we quantify the efficiency of wash-in leptogenesis, in relation to the optimal outcome mentioned above, $q_{B-L}=-\nicefrac{{3}}{{10}}\,q_{e}$, in terms of the efficiency factor
@@
In the following, we are therefore interested in identifying the region of parameter space where $\kappa\gtrsim 2.6\times 10^{-4}$. To this end, we perform a numerical scan over the four-dimensional parameter space spanned by $\widetilde{m}_{1}$, $M$, $p_{e}$, and $p_{\mu}$. More precisely, at each point in the $\widetilde{m}_{1}$ – $M$ plane, we scan over $p_{e}\in[0,1]$ and $p_{\mu}\in[0,1-p_{e}]$ and identify the branching ratios that either maximize $\kappa$ or $-\kappa$. This analysis provides us with two extremal solutions for the efficiency factor, $\kappa_{+}>0$ and $\kappa_{-}<0$, in the $\widetilde{m}_{1}$ – $M$ plane; see Fig. 2. Both solutions are physically viable, as we can always choose the sign of the initial $q_{e}^{\rm ini}$ charge appropriately.
@@
Conclusions — The low RHN mass bounds in Eq. (27) are a remarkable result, which highlights that, assuming a large primordial input charge near the CPI threshold, the viable RHN mass range for successful wash-in leptogenesis extends around one order of magnitude below the equilibration temperature of the electron Yukawa interaction, $M_{\rm min}\sim 0.1\,T_{e}$. Our new bound on the RHN mass scale is thus one order of magnitude lower than the naive value of $M_{\rm min}\sim 100\,\textrm{TeV}$ initially stated in Ref. Domcke:2020quw and roughly five orders of magnitude below the classic Davidson–Ibarra bound on the RHN mass in standard thermal leptogenesis. We thus arrive at our main conclusion, which we already anticipated in the introduction: if, in the context of baryogenesis via leptogenesis, RHN interactions only serve the purpose to violate $B\!-\!L$ and $CP$ violation is attributed to new chargegenesis dynamics at higher energies, large viable regions in the neutrino mass parameter space open up; see Fig. 2. The relaxation of the lower bound on the RHN mass scale $M_{N}$ in wash-in leptogenesis has notably important consequences for model building, as $M_{N}$ no longer needs to lie several orders of magnitude above the electroweak scale.

### v2

@@ Abstract
Leptogenesis is an attractive scenario for the generation of the baryon asymmetry of the Universe that relies on the dynamics of right-handed neutrinos (RHNs) in the seesaw extension of the Standard Model. In standard thermal leptogenesis, the RHN mass scale $M_{N}$ is subject to the Davidson–Ibarra bound, $M_{N}\gtrsim 10^{9}\,\textrm{GeV}$, which builds on the assumption that RHN decays are responsible for the violation of both charge-parity invariance ($CP$) and baryon-minus-lepton number (BLB\!-\!L). In this paper, we relax this assumption in the context of the more flexible framework of wash-in leptogenesis, in which $CP$ violation is encoded in the initial conditions and the only remaining task of the RHN decays is to violate BLB\!-\!L. Solving the relevant set of Boltzmann equations for vanishing initial baryon and lepton numbers (i.e., $B=L_{e}=L_{\mu}=L_{\tau}=0$ initially), we find that, in wash-in leptogenesis, the RHN mass scale can be as low as 7 TeV. Wash-in leptogenesis at such low RHN masses requires the presence of a primordial charge asymmetry between right-handed electrons and left-handed positrons. We discuss several possibilities for the origin of such an asymmetry and comment on its implications for the chiral instability of the Standard Model plasma.
preprint: MITP-25-002preprint: MS-TP-25-01
@@
Introduction — The baryon asymmetry of the Universe (BAU), commonly quantified in terms of the baryon-to-photon ratio $\eta_{B}^{\rm obs}=n_{B}^{0}/n_{\gamma}^{0}\simeq 6.12\times 10^{-10}$ Planck:2018vyg; ParticleDataGroup:2024cfk, cannot be explained by the Standard Model (SM) of particle physics and hence provides evidence for new physics beyond the SM (BSM) Sakharov:1967dj. An attractive BSM scenario capable of dynamically generating the BAU in the early Universe consists of baryogenesis via leptogenesis Fukugita:1986hr, which relies on the interactions of right-handed neutrinos (RHNs) in the seesaw extension of the SM Minkowski:1977sc; Yanagida:1979as; Yanagida:1980xy; Gell-Mann:1979vob; Mohapatra:1979ia, thus linking the generation of the BAU to the origin of the tiny masses of the left-handed neutrinos in the SM Buchmuller:2005eh; Chun:2017spz; Bodeker:2020ghk. In its standard implementation, leptogenesis proceeds via the out-of-equilibrium decays of heavy RHNs, which simultaneously violate charge–parity invariance ($CP$) and baryon-minus-lepton number (BLB\!-\!L). These decays first produce a primordial lepton asymmetry, which is subsequently converted by means of the chemical transport in the thermal plasma, including electroweak sphalerons Kuzmin:1985mm, to a primordial baryon asymmetry.
@@
The requirement of sufficient $CP$ violation in RHN decays bounds the RHN mass scale from below, $M_{N}\gtrsim 10^{9}\,\textrm{GeV}$. This constraint is known as the Davidson–Ibarra bound Davidson:2002qv; Buchmuller:2002rq (see Ref. Garbrecht:2024xfs for a recent refinement). It assumes a hierarchical RHN mass spectrum (e.g., $M_{3}\gg M_{2}\gg M_{1}\equiv M_{N}$ in models with three RHNs) and can be circumvented in the case of a quasi-degenerate RHN mass spectrum (e.g., $M_{3}\gg M_{2}\simeq M_{1}$ or $M_{3}\simeq M_{2}\simeq M_{1}$), which gives rise to the scenario of resonant leptogenesis Pilaftsis:2003gt; Pilaftsis:2005rv (see also Refs. Klaric:2020phc; Klaric:2021cpi). In this paper, we will, however, ignore the possibility of resonantly enhanced $CP$ asymmetry in RHN decays and focus instead on an alternative route to low-scale leptogenesis at RHN masses $M_{N}\ll 10^{9}\,\textrm{GeV}$: wash-in leptogenesis Domcke:2020quw. As we will show in this paper, wash-in leptogenesis can still successfully operate at RHN masses slightly below $10\,\textrm{TeV}$. In alternative scenarios, such as resonant leptogenesis or freeze-in leptogenesis via RHN oscillations (Akhmedov–Rubakov–Smirnov or ARS leptogenesis Akhmedov:1998qx), it is possible to produce the baryon asymmetry at even lower RHN masses, partially even in the sub-GeV regime Klaric:2021cpi. We emphasize that wash-in leptogenesis differs from these scenarios: unlike resonant leptogenesis, it requires no approximate mass degeneracy among the RHN mass states; and unlike both resonant and ARS leptogenesis, it does not depend on the amount of $CP$ violation in the RHN sector [see the discussion below Eq. (1) for more details].
@@
Indeed, the key idea behind wash-in leptogenesis is to liberate the RHN decays from the requirement of sufficient $CP$ violation. In wash-in leptogenesis, RHN interactions are treated as mere spectator processes and on the same footing as the electroweak sphalerons. From the perspective of wash-in leptogenesis, sphalerons and RHN decays represent nothing but sources of B+LB\!+\!L violation and BLB\!-\!L violation, respectively. $CP$ violation, on the other hand, is attributed to new $CP$-violating dynamics at high temperatures, $T\gg M_{N}$, which set the required $CP$-violating initial conditions for wash-in leptogenesis. In this sense, wash-in leptogenesis does not represent a complete baryogenesis scenario on its own, but calls for an ultraviolet completion — a chargegenesis mechanism that is responsible for the generation of primordial $CP$-violating charge asymmetries at high energies. Wash-in leptogenesis then acts on the nontrivial configuration of chemical potentials induced by chargegenesis and reshuffles the primordial charge asymmetries in a BLB\!-\!L-violating manner. The SM notably offers more than ten conserved global charges at high temperatures that could be produced during chargegenesis Domcke:2020quw, which renders baryogenesis via leptogenesis via chargegenesis significantly more versatile than baryogenesis via leptogenesis. As a consequence, the range of possible chargegenesis scenarios is extremely broad; primordial charge asymmetries suitable for wash-in leptogenesis can, e.g., be produced during axion inflation Domcke:2022kfs, by evaporating primordial black holes Schmitz:2023pfy, in the decay of heavy Higgs doublets Mukaida:2024eqi, or via new nonminimal gravitational interactions Mojahed:2024yus.
@@
As the temperature of the thermal bath drops, the last global charge to be erased by the SM interactions is the asymmetry between right-handed electrons and left-handed positrons, $q_{e}=n_{e}-\bar{n}_{e}$. Therefore, as long as the temperature scale of wash-in leptogenesis exceeds the equilibration temperature of the electron Yukawa interaction, $T_{e}\simeq 85\,\textrm{TeV}$ Bodeker:2019ajh, it is always possible to realize successful wash-in leptogenesis by assuming a nonzero primordial electron charge, $q_{e}^{\rm ini}\neq 0$. Correspondingly, the authors of Ref. Domcke:2020quw concluded that wash-in leptogenesis can take place at temperatures as low as $T\simeq 100\,\textrm{TeV}$. The purpose of the present paper is to make this statement more precise and study wash-in leptogenesis at $T\lesssim T_{e}$, where $q_{e}/T^{3}$ can no longer be treated as a constant, i.e., in a temperature regime where the asymmetry wash-in by RHN processes competes with the wash-out of the primordial $q_{e}$ input charge due to the progressing equilibration of the electron Yukawa interaction. To this end, we shall extend the treatment in Ref. Domcke:2020quw and supplement the Boltzmann equations for the three flavor asymmetries $q_{\Delta_{\alpha}}$, where $\Delta_{\alpha}=B/3-L_{\alpha}$ and $\alpha=e,\mu,\tau$, that are being produced during wash-in leptogenesis by a Boltzmann equation for the time-dependent electron charge $q_{e}$. Solving this coupled set of Boltzmann equations will then allow us to conclude that wash-in leptogenesis is able to produce the BAU for RHN masses as low as $M_{N}\simeq 7\,\textrm{TeV}$. In summary, in standard thermal leptogenesis where RHN decays act as a source of both $CP$ and BLB\!-\!L violation, the RHN mass scale must satisfy the Davidson–Ibarra bound, $M_{N}\gtrsim 10^{9}\,\textrm{GeV}$; if RHN decays only act as a source of BLB\!-\!L violation, it suffices to demand that $M_{N}\gtrsim 7\,\textrm{TeV}$.
@@
$\left(\partial_{t}+3H\right)q_{\Delta_{\alpha}}=\varepsilon_{1\alpha}\,\Gamma_{1}\left(n_{N_{1}}-n_{N_{1}}^{\rm eq}\right)-\sum_{\beta}\gamma_{\alpha\beta}^{\rm w}\,\frac{\mu_{\ell_{\beta}}+\mu_{\phi}}{T}\,.$ (1)
Here, the negative sign on the left-hand side reflects the fact that the lepton flavor number $L_{\alpha}$ enters $\Delta_{\alpha}$ with a negative sign. The two terms on the right-hand side correspond to the standard source term for $q_{\Delta_{\alpha}}$ and the standard wash-out term in thermal leptogenesis, respectively. Here, the total wash-out interaction density receives contributions from RHN inverse decays as well as from lepton-number violating ($\Delta L=2$) and lepton-flavor-violating ($\Delta L=0$) processes. This interaction density is multiplied by $\mu_{\ell_{\beta}}+\mu_{\phi}$, i.e., the sum of the chemical potentials of the SM charged-lepton and Higgs doublets. At $T\sim T_{e}$, the chemical potentials of the 16 SM particle multiplets, $\mu_{i}$ ($i=e,\mu,\tau,\ell_{e},\ell_{\mu},\ell_{\tau},u,c,t,d,s,b,Q_{1},Q_{2},Q_{3},\phi$), are subject to 16 constraints: 11 conditions for 11 linearly independent Yukawa and sphaleron interactions in chemical equilibrium and 5 conservation laws for 5 conserved global charges ($\Delta_{e,\mu,\tau}$, hypercharge $Y$, and the asymmetry in right-handed electrons). Solving this linear system of constraint equations for the 16 chemical potentials $\mu_{i}$, one finds Domcke:2020kcp; Domcke:2020quw
@@
or in more compact notation,
$\mu_{\ell_{\alpha}}+\mu_{\phi}=\bar{\mu}_{\alpha}^{0}-\sum_{\beta}C_{\alpha\beta}\,\bar{\mu}_{\Delta_{\beta}}\,.$ (3)
where the second term on the right-hand side contains the standard flavor coupling matrix $C_{\alpha\beta}$ Barbieri:1999ma; Abada:2006fw; Nardi:2006fx; Abada:2006ea; Blanchet:2006be; Antusch:2006cw induced by the spectator processes in the thermal bath Buchmuller:2001sr; Garbrecht:2014kda; Garbrecht:2019zaa, while the first term is the relevant term for wash-in leptogenesis.
Inserting the affine relation in Eq. (3) into Eq. (1), the Boltzmann equation for the three lepton flavor asymmetries can be brought into the following form,
$-\left(\partial_{t}+3H\right)q_{\Delta_{\alpha}}=$ (4)
$\varepsilon_{1\alpha}\,\Gamma_{1}\left(n_{N_{1}}-n_{N_{1}}^{\rm eq}\right)-\sum_{\beta}\Gamma_{\alpha\beta}^{\rm w}\left(q_{\beta}^{0}-\sum_{\sigma}C_{\beta\sigma}\,q_{\Delta_{\sigma}}\right)\,$,
where $q_{\alpha}^{0}\equiv\nicefrac{{1}}{{6}}\>\bar{\mu}_{\alpha}^{0}T^{2}$, $q_{\Delta_{\alpha}}\equiv\nicefrac{{1}}{{6}}\>\bar{\mu}_{\Delta_{\alpha}}T^{2}$, and $\Gamma_{\alpha\beta}=6/T^{3}\,\gamma_{\alpha\beta}$. This evolution equation is linear in the $q_{\Delta_{\alpha}}$, which allows us to write $q_{\Delta_{\alpha}}=q_{\Delta_{\alpha}}^{\rm th}+q_{\Delta_{\alpha}}^{\rm win}$ and split Eq. (4) into two equations,
$-\left(\partial_{t}+3H\right)q_{\Delta_{\alpha}}^{\rm th}$ $=\varepsilon_{1\alpha}\,\Gamma_{1}\left(n_{N_{1}}-n_{N_{1}}^{\rm eq}\right)-\sum_{\beta,\sigma}\Gamma_{\alpha\beta}^{\rm w}C_{\beta\sigma}\,q_{\Delta_{\sigma}}^{\rm th}\,$, (5)
$-\left(\partial_{t}+3H\right)q_{\Delta_{\alpha}}^{\rm win}$ $=-\sum_{\beta}\Gamma_{\alpha\beta}^{\rm w}\left(q_{\beta}^{0}-\sum_{\sigma}C_{\beta\sigma}\,q_{\Delta_{\sigma}}^{\rm win}\right)\,.$ (6)
Indeed, the sum of these two equations returns again the original Boltzmann equation in Eq. (4). The first equation, Eq. (5), is identical to the standard Boltzmann equation for thermal leptogenesis (possibly in the resonant regime) in a trivial chemical background (i.e., no affine $q_{\alpha}^{0}$ term) and fully accounts for the effect of $CP$ violation in the RHN sector. The second equation, Eq. (6), on the other hand, describes the evolution of $q_{\Delta_{\alpha}}^{\rm win}$, i.e., the contributions to the lepton flavor asymmetries produced during wash-in leptogenesis. By definition, Eq. (6) is completely independent of the RHN $CP$ violation parameters $\varepsilon_{1\alpha}$. In the following, we shall focus on the wash-in contributions to the total lepton flavor asymmetries and assume that $q_{\Delta_{\alpha}}^{\rm th}\ll q_{\Delta_{\alpha}}^{\rm win}$ for all $\alpha$. In other words, we shall work in a part of the RHN parameter space where the RHN $CP$ violation parameters $\varepsilon_{1\alpha}$ are negligibly small and neither resonant nor ARS leptogenesis manage to explain the observed value of the BAU. Instead, we will assume a nonvanishing primordial charge asymmetry, $\bar{\mu}_{e}\neq 0$, which significantly changes the character of the processes contributing to the rate $\Gamma_{\alpha\beta}^{\rm w}$ in Eq. (6): Typically, these processes describe wash-out effects that drive the $q_{\Delta_{\alpha}}$ asymmetries to zero; see Eq. (5). However, in the presence of a primordial input charge $q_{e}$, the wash-out processes begin to wash in nonzero lepton flavor asymmetries proportional to $q_{e}$, even if $q_{\Delta_{\alpha}}=0$ initially. We notably attribute the origin of the primordial charge asymmetry $q_{e}$ to new $CP$-violating dynamics in a different sector (see the discussion above), such that every single term in Eq. (6) is insensitive to the effect of $CP$ violation in the RHN sector.
Assuming a constant comoving charge density, $q_{e}/T^{3}=\textrm{const}$, wash-in leptogenesis in the strong wash-in regime (conventionally, strong wash-out regime) at $T\gtrsim T_{e}$ results in a BLB\!-\!L asymmetry $q_{B-L}=-\nicefrac{{3}}{{10}}\,q_{e}$ Domcke:2020quw. Around $T\sim T_{e}$, the assumption of constant $q_{e}/T^{3}$, however, begins to break down and $q_{e}/T^{3}$ needs to be described as a time-dependent quantity. The relevant Boltzmann equation for $q_{e}$ around the equilibration temperature of the electron Yukawa interaction has been derived in Ref. Bodeker:2019ajh,
@@
Meanwhile, the electron equilibration rate can be written as $\Gamma_{e}=C_{e}\,T$, where the coefficient $C_{e}$ receives contributions from $2\rightarrow 2$ processes involving one electron Yukawa vertex as well as from $1\leftrightarrow 2$ decays and inverse decays of the SM Higgs bosons, right-handed electrons and left-handed leptons. In fact, the latter processes also receive corrections from $1n\leftrightarrow 2n$ scatterings with soft gauge boson exchanges, which can be taken care of by Landau–Pomeranchuk–Migdal (LPM) resummation Landau:1953um; Landau:1953gr; Migdal:1956tc. Taking all these processes into account, the authors of Ref. Bodeker:2019ajh find
@@
where $_{t}, cc_{\ell}$, and $c_{e}$ receive contributions from both $2\rightarrow 2$ processes and LPM-resummed $1n\leftrightarrow 2n$ processes,
@@
These initial conditions define the minimal scenario that we are interested in in this paper, in which the RHN interactions during wash-in leptogenesis are the only source of lepton number and lepton flavor violation. In particular, the first condition corresponds to the standard starting point of other typical leptogenesis scenarios in the literature: before the dynamics in the RHN sector become relevant, we set $Q_{\Delta_{\alpha}}=0$ for all $\alpha$. This assumption imposes restrictions on the chargegenesis mechanism setting the initial conditions for wash-in leptogenesis. Our analysis in this paper only applies to scenarios of chargegenesis that result in $Q_{e}\neq 0$, possibly other charges, but $Q_{\Delta_{\alpha}}=0$ for all $\alpha$. Chargegenesis may, e.g., also lead to nonzero $Q_{\mu}\neq Q_{e}$ and nonzero $Q_{\tau}\neq Q_{e}$, meaning that flavor universality may be violated during chargegenesis. Any nonzero primordial muon or tau charge asymmetries will simply be washed out by the respective Yukawa interactions before the temperature regime relevant for our analysis in this paper is reached. This is explicitly illustrated by Eqs. (4.6) and (4.8)–(4.11) in Ref. Domcke:2022kfs, which are general and go beyond the specific scenario discussed in Ref. Domcke:2022kfs. Our analysis in this paper applies to any chargegenesis scenario that populates the entries of the charge vectors on the right-hand side of these equations, except for the lepton flavor asymmetries $Q_{\Delta_{\alpha}}$.
Strictly speaking, the framework of wash-in leptogenesis is even more general and does not require vanishing initial conditions for the lepton flavor asymmetries. Similarly as in the case of other leptogenesis scenarios in the literature, our analysis may be extended by allowing for the possibility of pre-existing lepton flavor asymmetries; see, e.g., Ref. Bertuzzo:2010et, which discusses the impact of pre-existing asymmetries on flavored thermal leptogenesis. In the strong wash-in regime, the effect of nonvanishing initial conditions, $Q_{\Delta_{\alpha}}\left(z\ll 1\right)\neq 0$, will be exponentially suppressed, as illustrated by the exact analytical solution in Eq. (7) in Ref. Domcke:2020quw. Nonvanishing initial conditions in the weak wash-in regime, however, deserve more attention in future work.
Four example solutions of our Boltzmann equations for these initial conditions are shown in Fig. 1. For a given numerical solution, we quantify the efficiency of wash-in leptogenesis, in relation to the optimal outcome mentioned above, $q_{B-L}=-\nicefrac{{3}}{{10}}\,q_{e}$, in terms of the efficiency factor
@@
In the following, we are therefore interested in identifying the region of parameter space where $\kappa\gtrsim 2.6\times 10^{-4}$. To this end, we perform a numerical scan over the four-dimensional parameter space spanned by $\widetilde{m}_{1}$, $M$, $p_{e}$, and $p_{\mu}$. More precisely, at each point in the $\widetilde{m}_{1}$ – $M$ plane, we scan over $p_{e}\in[0,1]$ and $p_{\mu}\in[0,1-p_{e}]$ and identify the branching ratios that either maximize $\kappa$ or $\kappa$. This analysis provides us with two extremal solutions for the efficiency factor, $\kappa_{+}>0$ and $\kappa_{-}<0$, in the $\widetilde{m}_{1}$ – $M$ plane; see Fig. 2. Both solutions are physically viable, as we can always choose the sign of the initial $q_{e}^{\rm ini}$ charge appropriately.
@@
Conclusions — The low RHN mass bounds in Eq. (31) are a remarkable result, which highlights that, assuming a large primordial input charge near the CPI threshold, the viable RHN mass range for successful wash-in leptogenesis extends around one order of magnitude below the equilibration temperature of the electron Yukawa interaction, $M_{\rm min}\sim 0.1\,T_{e}$. Our new bound on the RHN mass scale is thus one order of magnitude lower than the naive value of $M_{\rm min}\sim 100\,\textrm{TeV}$ initially stated in Ref. Domcke:2020quw and roughly five orders of magnitude below the classic Davidson–Ibarra bound on the RHN mass in standard thermal leptogenesis. We thus arrive at our main conclusion, which we already anticipated in the introduction: if, in the context of baryogenesis via leptogenesis, RHN interactions only serve the purpose to violate BLB\!-\!L and $CP$ violation is attributed to new chargegenesis dynamics at higher energies, large viable regions in the neutrino mass parameter space open up; see Fig. 2. The relaxation of the lower bound on the RHN mass scale $M_{N}$ in wash-in leptogenesis has notably important consequences for model building, as $M_{N}$ no longer needs to lie several orders of magnitude above the electroweak scale.